# Thermodynamics of Modified Chaplygin Jacobi And Modified Chaplygin Abel Gas: Stability Analysis and Observational Constraints

### v1

@@ Abstract
This paper explores the thermodynamic properties and stability of two newly introduced gas models, namely the Modified Chaplygin Jacobi gas and the Modified Chaplygin Abel gas. To achieve this, we examine the behavior of relevant physical parameters to gain in depth information about the evolution of the universe. The specific heat formalism is employed to verify the applicability of the third law of thermodynamics. Furthermore, the equation of state for the thermal system is obtained by applying thermodynamic variables. The stability of the gas models is investigated within the framework of classical thermodynamics, focusing on adiabatic processes, specific heat capacities, and isothermal conditions. It is inferred that the proposed fluid configurations exhibit thermodynamic stability and undergo adiabatic expansion for suitable parameter choices. We then perform observational analysis using CC+BAO and Pantheon+SH0ES datasets to impose constraints on our model parameters using the Markov Chain Monte Carlo (MCMC) process.
@@ 1   Introduction
The present cosmological observations from the Type Ia supernovae indicate our universe to be expanding in an accelerated manner [1, 2, 3] and hence contradict the strong energy condition, which implies $\rho+3p<0$. The exotic matter that drives the universe to expand acceleratedly is known to be “Dark Energy” [4, 5]. The straightforward approach to resolve this problem was to take the famous cosmological constant $\Lambda$ as a suitable choice for dark energy taking the equation of state (EoS) parameter $\omega=\frac{p}{\rho}=-1$ as constant. However this approach has faced several fatal drawbacks that remain unanswered [6, 7, 8]. To overcome this problem, several dark energy models with dynamic vacuum energy have been proposed. This theory introduces an additional degree of freedom, characterized by a scalar field. The Lagrangian incorporates a potential or kinetic term, which facilitates the transition of the universe from a dust-like matter-dominated state to one experiencing accelerated expansion. The quintessence model [9, 10, 11, 12] is one that kind that is mostly favored by physicists to be a potential candidate for dark energy. Additionally, alternative dark energy models featuring dynamic vacuum energy have been explored, including tachyon fields [13], k-essence [14], H-essence [15], dilaton [16], phantom fields [17], DBI-quintessence [18], and DBI-essence [19]. The Chaplygin gas model, initially introduced in the context of aerodynamics by Sergey A. Chaplygin in 1904 [20],has recently attracted considerable interest in both cosmology and theoretical physics because of its distinctive equation of state (EoS),
$p=-\dfrac{B}{\rho}$ (1)
@@
Where, $B$ is a positive parameter. The consideration of using Chaplygin gas in a FRW universe for the explanation of the accelerated expansion of the universe was first proposed by Kamenshchik et al. [21, 22]. Equation (1) reveals a fascinating link to string theory, as it originates from the Nambu-Goto action, which governs the dynamics of D-branes moving within a $(d+2)$-dimensional spacetime under the light-cone gauge  [23]. In addition, the Chaplygin gas is the only known fluid that allows for a supersymmetric extension  [24, 25]. The Chaplygin gas-like equation of state can then be derived from a tachyon scalar field model, which is described by a Dirac-Born-Infeld type Lagrangian with a constant potential for the field [26, 27]. A significant aspect of the Chaplygin gas model is its ability to unify dark matter and dark energy within a single theoretical framework [28, 29]. But, this model needed to explain the structure formation era in the universe [30, 31]. Bento et al. [32] introduced a more general version of the Chaplygin gas EoS and it is referred to as Generalised Chaplygin gas (GCG). Another modification to the equation of state was introduced and is known as Modified Chaplygin Gas(MCG)[33, 34]. This MCG model can reproduce radiation dominated era for $A=\frac{1}{3}$ for minuscule values of scale factor. The MCG model has limitations in accurately describing the universe’s behavior between the early and late acceleration periods [35]. To improve the agreement between models and observational data, various limits on the model parameters of both the Chaplygin Gas (CG) and Modified Chaplygin Gas (MCG) were explored [36, 37, 38, 39, 40, 41]. Villanueva proposed the generalized Chaplygin-Jacobi gas through the application of Jacobi’s elliptic function  [42, 43], which is linked to the generalized Chaplygin scalar field. This model is capable of describing the inflationary phase and has been tested against the data recently provided by the Planck 2015 mission. Recently, Debnath [44] introduced two models of Chaplygin gas namely, Modified Chaplygin Jacobi Gas(MCJG) and Modified Chaplygin Abel Gas(MCAG). He replaced the hyperbolic cosine function in the generating function of the scalar field with the Jacobi and Abel elliptic functions, deriving the equation that connects the pressure and density of the MCJG and MCAG.
@@
The intersection of black hole thermodynamics and Einstein’s field equations emerged as a fundamental concept within Einstein gravity framework [45]. This connection established that the entropy-area proportionality principle, coupled with the fundamental thermodynamic relation $\delta Q=TdS$ in Rindler spacetime, leads naturally to Einstein’s equations. The correspondence between thermodynamic and geometric quantities in black hole physics reveals itself through entropy’s relationship to horizon area and temperature’s connection to surface gravity [46]. A significant advancement occurred when Verlinde demonstrated [47] that the Friedmann equation in a radiation-dominated FRW universe exhibits structural similarities to the Cardy-Verlinde formula, which describes entropy relationships in conformal field theory. The thermodynamic properties of de Sitter spacetime were first elucidated through the pioneering work of Gibbons and Hawking [48]. Further developments by Gong and colleagues [49] established the apparent horizon as the most relevant boundary for thermodynamic analysis, based on the established relationship between Friedmann equations and thermodynamic principles. This becomes particularly significant in accelerating universes dominated by dark energy, where the event horizon differs from the apparent horizon. Research has shown that thermodynamic laws maintain validity at the apparent horizon but fail at the event horizon when these boundaries are distinct [50]. Investigations into Chaplygin gas-dominated universes [51] examined the generalized second law (GSL) in relation to event horizons in accelerating cosmological models, confirming GSL validity during early expansion phases. A comprehensive comparative study [52] evaluated GSL validity at both cosmological horizons (apparent and event) in FRW universes containing various Chaplygin gas models. The thermodynamic framework has been extensively explored across multiple variations:
@@
This extensive body of research provides the foundation for our investigation into the thermodynamic characteristics of Modified Chaplygin Jacobi and Modified Chaplygin Abel gas parameters.
The paper aims to analyse the thermodynamic behaviour of Modified Chaplygin Jacobi(MCJG) and Modified Chaplygin Abel gas(MCAG) and also study their temperature behaviour. Also the thermodynamic stability of MCJG and MCAG using thermal and adiabatic equation of states is part of our analysis in this work. The paper is arranged as follows: In Sect. 2 and 3 we have approximate MCJG and MCAG model and analysed their thermodynamic behaviour and stability, Sect. 4 is dedicated to study the model parameters and testing the viability of our approximations using CC+BAO and Pantheon+SH0ES datasets and finally we wind up our paper in Sect. 5 with concluding remarks.
2   Thermodynamic Analysis of Modified Chaplygin Jacobi Gas
@@
$p=\left[(2K-1)(1+A)-1\right]\rho-\frac{KB}{\rho^{\alpha}}+\frac{(1-K)(1+A)^{2}\rho^{2+\alpha}}{B}$ (2)
In this context, $A$ and $B$ are positive constant parameters, with $0\leq\alpha\leq 1$ and $0\leq K\leq 1$ representing the modulus of the elliptic function. This expression defines the pressure-density relationship of the “Modified Chaplygin Jacobi Gas (MCJG)”.
From the general Thermodynamics we know that[58],
$\rho=\frac{U}{V}$ (3)
@@
$p=-\left(\frac{\partial U}{\partial V}\right)_{S}$ (4)
@@
$U=V\left(\dfrac{B\left(1+\dfrac{1}{-1+K-e^{(1+A)(1+\alpha)BC_{1}}V^{(A+1)(\alpha+1)}}\right)}{1+A}\right)^{\frac{1}{1+\alpha}}$ (5)
@@
$\rho=\left[\frac{B}{1+A}\left(1-\dfrac{1}{1-K-(\frac{B}{C})^{(A+1)}V^{(A+1)(\alpha+1)}}\right)\right]^{\frac{1}{1+\alpha}}$ (7)
@@
$\rho=\left[\frac{B}{A+1}\left(1+2K\left(\dfrac{V_{0}}{V}\right)^{(\alpha+1)(A+1)}\right)\right]^{\dfrac{1}{1+\alpha}}$ (8)
@@
$\rho\approx\left[\left(\dfrac{B}{A+1}\right)^{\frac{1}{1+\alpha}}+\dfrac{2K}{\alpha+1}\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}\right]$ (10)
@@
$p=-BK\left[\frac{B}{A+1}\left(1+\dfrac{\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}}{1+K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}}\right)\right]^{-\dfrac{\alpha}{\alpha+1}}$ (11)
@@
$p$ $\approx-BK\left[\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}+\left(1+\dfrac{1}{K(\alpha+1)}\right)\right]^{-\alpha}$ (12)
$\approx-K\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}+K\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}$ (13)
$\approx 0$ (14)
@@
$p\approx-K\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}$ (15)
@@
Figure 1: A plot showing the dark energy density $\rho$ and pressure $p$ as functions of volume $V$ for the Modified Chaplygin Jacobi Gas (MCJG) model.
@@ 2.1   EoS Parameter
$\omega=\dfrac{p}{\rho}$
. From the equations (11) and (8) we have obtained,
$\omega=-\dfrac{BK}{\left[\dfrac{B}{A+1}\left(2K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)\right]^{\frac{1}{\alpha+1}}\left[\dfrac{B\left(K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}{(A+1)\left(K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}\right]^{\frac{\alpha}{\alpha+1}}}$ (16)
It represents a very small constant value of EoS parameter at small volume ($V\to 0$) and at large volume ($V\to\infty$) the EoS parameter becomes negative and volume dependent and is plotted in fig. 2(a).

### v2

# Thermodynamics of Modified Chaplygin-Jacobi Gas and Modified Chaplygin-Abel Gas: Stability Analysis and Observational Constraints

@@ Abstract
This paper explores the thermodynamic properties and stability of two newly introduced gas models, namely the Modified Chaplygin-Jacobi gas and the Modified Chaplygin-Abel gas. To achieve this, we examine the behavior of relevant physical parameters to gain in depth information about the evolution of the universe. The specific heat formalism is employed to verify the applicability of the third law of thermodynamics. Furthermore, the equation of state for the thermal system is obtained by applying thermodynamic variables. The stability of the gas models is investigated within the framework of classical thermodynamics, focusing on adiabatic processes, specific heat capacities, and isothermal conditions. It is inferred that the proposed fluid configurations exhibit thermodynamic stability and undergo adiabatic expansion for suitable parameter choices. We then perform observational analysis using CC+BAO and Pantheon+SH0ES datasets to impose constraints on our model parameters using the Markov Chain Monte Carlo (MCMC) process.
@@ 1   Introduction
The present cosmological observations from the Type Ia supernovae indicate our universe to be expanding in an accelerated manner [1, 2, 3] and hence contradict the strong energy condition, which implies $\rho+3p<0$. The exotic matter that drives the universe to expand acceleratedly is known to be “Dark Energy” [4, 5]. The straightforward approach to resolve this problem was to take the famous cosmological constant $\Lambda$ as a suitable choice for dark energy taking the equation of state (EoS) parameter $\omega=\frac{p}{\rho}=-1$ as constant. However this approach has faced several fatal drawbacks that remain unanswered [6, 7, 8]. To overcome this problem, several dark energy models with dynamic vacuum energy have been proposed. This theory introduces an additional degree of freedom, characterized by a scalar field. The Lagrangian incorporates a potential or kinetic term, which facilitates the transition of the universe from a dust-like matter-dominated state to one experiencing accelerated expansion. The quintessence model [9, 10, 11, 12] is one that kind that is mostly favored by physicists to be a potential candidate for dark energy. Additionally, alternative dark energy models featuring dynamic vacuum energy have been explored, including tachyon fields [13], k-essence [14], H-essence [15], dilaton [16], phantom fields [17], DBI-quintessence [18], and DBI-essence [19]. The Chaplygin gas model, initially introduced in the context of aerodynamics by Sergey A. Chaplygin in 1904 [20], has recently attracted considerable interest in both cosmology and theoretical physics because of its distinctive equation of state (EoS). The consideration of using Chaplygin gas in a FRW universe for the explanation of the accelerated expansion of the universe was first proposed by Kamenshchik et al. [21, 22]. Chaplygin gas equation of state reveals a fascinating link to string theory, as it originates from the Nambu-Goto action, which governs the dynamics of D-branes moving within a $(d+2)$-dimensional spacetime under the light-cone gauge  [23]. In addition, the Chaplygin gas is the only known fluid that allows for a supersymmetric extension  [24, 25]. The Chaplygin gas-like equation of state can then be derived from a tachyon scalar field model, which is described by a Dirac-Born-Infeld type Lagrangian with a constant potential for the field [26, 27]. A significant aspect of the Chaplygin gas model is its ability to unify dark matter and dark energy within a single theoretical framework [28, 29]. But, this model needed to explain the structure formation era in the universe [30, 31]. Bento et al. [32] introduced a more general version of the Chaplygin gas EoS and it is referred to as Generalised Chaplygin gas (GCG). Another modification to the equation of state was introduced and is known as Modified Chaplygin Gas(MCG)[33, 34]. The MCG model has limitations in accurately describing the universe’s behavior between the early and late acceleration periods [35]. To improve the agreement between models and observational data, various limits on the model parameters of both the Chaplygin Gas (CG) and Modified Chaplygin Gas (MCG) were explored [36, 37, 38, 39, 40, 41]. Villanueva proposed the generalized Chaplygin-Jacobi gas through the application of Jacobi’s elliptic function  [42, 43], which is linked to the generalized Chaplygin scalar field. This model is capable of describing the inflationary phase and has been tested against the data recently provided by the Planck 2015 mission. Recently, Debnath [44] introduced two models of Chaplygin gas namely, Modified Chaplygin-Jacobi Gas (MCJG) and Modified Chaplygin-Abel Gas (MCAG). He replaced the hyperbolic cosine function in the generating function of the scalar field with the Jacobi and Abel elliptic functions, deriving the equation that connects the pressure and density of the MCJG and MCAG. The reason behind using such elliptic functions is as follows: The exact solution of the Einstein field equations with an inflation field is very much hard to achieve. Still using slow roll approximations and other methods several literature suggests that the inflationary potentials has the form of trigonometric and hyperbolic functions [45, 46, 47, 48, 49, 50]. The elliptic functions are more generalised functions which reduce to trigonometric and hyperbolic forms as a special case. So, the use of elliptic functions in the inflationary potential and then the further extension of this function to obtain Modified Chaplygin gas like equation of state provide more generalisation, from both mathematical and physical perspectives.
@@
The intersection of black hole thermodynamics and Einstein’s field equations emerged as a fundamental concept within Einstein gravity framework [51]. This connection established that the entropy-area proportionality principle, coupled with the fundamental thermodynamic relation $\delta Q=TdS$ in Rindler spacetime, leads naturally to Einstein’s equations. The correspondence between thermodynamic and geometric quantities in black hole physics reveals itself through entropy’s relationship to horizon area and temperature’s connection to surface gravity [52]. A significant advancement occurred when Verlinde demonstrated [53] that the Friedmann equation in a radiation-dominated FRW universe exhibits structural similarities to the Cardy-Verlinde formula, which describes entropy relationships in conformal field theory. The thermodynamic properties of de Sitter spacetime were first elucidated through the pioneering work of Gibbons and Hawking [54]. Further developments by Gong and colleagues [55] established the apparent horizon as the most relevant boundary for thermodynamic analysis, based on the established relationship between Friedmann equations and thermodynamic principles. This becomes particularly significant in accelerating universes dominated by dark energy, where the event horizon differs from the apparent horizon. Research has shown that thermodynamic laws maintain validity at the apparent horizon but fail at the event horizon when these boundaries are distinct [56]. Recently, another work has showed that the thermodynamics of any matter fields including the Chaplygin gas is irreversible in nature, when satisfying the Landauer principle [57]. Investigations into Chaplygin gas-dominated universes [58] examined the generalized second law (GSL) in relation to event horizons in accelerating cosmological models, confirming GSL validity during early expansion phases. A comprehensive comparative study [59] evaluated GSL validity at both cosmological horizons (apparent and event) in FRW universes containing various Chaplygin gas models. The thermodynamic framework has been extensively explored across multiple variations:
@@
This extensive body of research provides the foundation for our investigation into the thermodynamic characteristics of MCJG and MCAG parameters.
The paper aims to analyse the thermodynamic behaviour of MCJG and MCAG and also study their temperature behaviour. Also the thermodynamic stability of MCJG and MCAG using thermal and adiabatic equation of states is part of our analysis in this work. The paper is arranged as follows: In Sect. 2 and 3 we have approximate MCJG and MCAG model and analysed their thermodynamic behaviour and stability, Sect. 4 is dedicated to study the model parameters and testing the viability of our approximations using CC+BAO and Pantheon+SH0ES datasets and finally we wind up our paper in Sect. 5 with concluding remarks.
2   Thermodynamic Analysis of Modified Chaplygin-Jacobi Gas
@@
$p=\left[(2K-1)(1+A)-1\right]\rho-\frac{KB}{\rho^{\alpha}}+\frac{(1-K)(1+A)^{2}\rho^{2+\alpha}}{B}\quad.$ (1)
In this context, $A$ and $B$ are positive constant parameters, with $0\leq\alpha\leq 1$ and $0\leq K\leq 1$ representing the modulus of the elliptic function. This expression defines the pressure-density relationship of the “Modified Chaplygin-Jacobi Gas”. The constant parameters $A$, $B$ and $\alpha$ are coming from the Modified Chaplygin gas and generalised Chaplygin gas equation of state [34, 32] and the detailed observational constraints of these parameters are given in ref [65, 66, 67, 68, 69]. Here, $B$ is a positive constant that influences the pressure term inversely proportional to the energy density. It plays a significant role in the thermodynamic stability and the transition between different cosmological eras [70, 71] [^5] In the aforementioned references this parameter is identified as $A$, but physically that does not make any difference as long as the form of the EoS of the MCG and GCG is preserved. Readers can tally the form of the EoS of MCG with ref. [34] to avoid any kind of confusion.. The value of $B$ affects the stability of the model and the collapse time rate in structure formation scenarios [72]. $A$ is a positive constant that modifies the linear term in the equation of state. It is crucial for describing the current accelerated expansion of the universe and is constrained by observational data such as CMBR measurements and The range of $A$ is typically constrained to ensure compatibility with observational data, with values often found between $-0.35$ and $0.025$ [70, 73] [^6] Here, also the same goes for the parameter $A$. The authors of the above references use $B$ instead of $A$ but physically they both signify the same thing.. For the values of $A=\frac{1}{3}$ the MCG model can reproduce radiation dominated era for minuscule values of scale factor. In some models, $A$ can be a function of the scale factor, adding variability to the equation of state and affecting the stability and thermodynamic properties of the gas [74]. The $\alpha$ parameter determines the non-linearity of the inverse density term. It ranges between 00 and $1$ and influences the transition from radiation-dominated to quintessence-dominated eras [71, 75]. The value of $\alpha$ affects the thermodynamic stability and the sound speed of the gas, with higher values leading to different stability conditions [76, 63] and it also plays a role in the evolution of perturbations and the turn-around redshifts in structure formation models [71]. The choice of the elliptic modulus $K=1$ reduced the generating function of MCJG to the generating function of MCG [44]. From the general Thermodynamics we know that[77],
$\rho=\frac{U}{V}\quad,$ (2)
@@
$p=-\left(\frac{\partial U}{\partial V}\right)_{S}.$ (3)
@@
$U=V\left(\dfrac{B\left(1+\dfrac{1}{-1+K-e^{(1+A)(1+\alpha)BC_{1}}V^{(A+1)(\alpha+1)}}\right)}{1+A}\right)^{\frac{1}{1+\alpha}}.$ (4)
@@
$\rho=\left[\frac{B}{1+A}\left(1-\dfrac{1}{1-K-(\frac{B}{C})^{(A+1)}V^{(A+1)(\alpha+1)}}\right)\right]^{\frac{1}{1+\alpha}}.$ (6)
@@
$\rho=\left[\frac{B}{A+1}\left(1+2K\left(\dfrac{V_{0}}{V}\right)^{(\alpha+1)(A+1)}\right)\right]^{\dfrac{1}{1+\alpha}}.$ (7)
@@
$\rho\approx\left[\left(\dfrac{B}{A+1}\right)^{\frac{1}{1+\alpha}}+\dfrac{2K}{\alpha+1}\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}\right].$ (9)
@@
$p=-BK\left[\frac{B}{A+1}\left(1+\dfrac{\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}}{1+K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}}\right)\right]^{-\dfrac{\alpha}{\alpha+1}}.$ (10)
@@
$p$ $\approx-BK\left[\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}+\left(1+\dfrac{1}{K(\alpha+1)}\right)\right]^{-\alpha}$
$\approx-K\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}+K\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}$
$\approx 0,$ (11)
@@
$p\approx-K\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}.$ (12)
@@
Figure 1: A plot showing the dark energy density $\rho$ and pressure $p$ as functions of volume $V$ for the Modified Chaplygin-Jacobi Gas model.
@@ 2.1   EoS Parameter
$\omega=\dfrac{p}{\rho}\quad.$
From the equations (10) and (7) we have obtained,
$\omega=-\dfrac{BK}{\left[\dfrac{B}{A+1}\left(2K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)\right]^{\frac{1}{\alpha+1}}\left[\dfrac{B\left(K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}{(A+1)\left(K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}\right]^{\frac{\alpha}{\alpha+1}}}\quad.$ (13)
It represents a very small constant value of EoS parameter at small volume ($V\to 0$) and at large volume ($V\to\infty$) the EoS parameter becomes negative and volume dependent and is plotted in fig. 2(a). It is interesting to show that for the value of parameters $A=V_{0}=0$, $B=K=1$ the EoS expressed in (13) reduces to $\omega=-1$ which corresponds to the original dark energy equation of state according to the $\Lambda$CDM model for any values of $\alpha$.